# Intention Knowledge Graph Construction for User Intention Relation Modeling

### v2

@@ Abstract
Understanding user intentions is challenging for online platforms. Recent work on intention knowledge graphs addresses this, but often lacks focus on connecting intentions, which is crucial for modeling user behavior and predicting future actions. This paper introduces a framework to automatically generate an intention knowledge graph, capturing connections between user intentions. Using the Amazon m2 dataset, we construct an intention graph with 351 million edges, demonstrating high plausibility and acceptance. Our model effectively predicts new session intentions and enhances product recommendations, outperforming previous state-of-the-art methods and showcasing the approach’s practical utility. [^1] We will release all data and code for this paper after this paper is accepted.
@@ 4.2 Automatic Evaluation
This study evaluates the task of linking unseen sessions to an existing intention graph. The dataset, comprising all session-intention edges, was split into training, validation, and test sets (8:1:1) based on sessions, with each session linked to 2-4 positive intentions. For evaluation, we applied random negative sampling to generate 30 candidate intentions per session, ranking the positive intentions among them. We compared large language models (Mistral-7B-Instruct-v0.1, Meta-Llama-3-8B-Instruct, and flan-t5-large) to our RIG model. Baselines use perplexity scores derived via proper prompting to rank intentions, while RIG employs an embedding-based approach. Specifically, we generate embeddings for sessions and intentions using a sentence model and use SASRec as a session encoder to compute session representations. As shown in Table 7, our RIG model achieves competitive accuracy (e.g., best Hit@10) while significantly outperforming LLMs in inference speed (3.01ms vs. 2,021–4,102ms), demonstrating its practicality for real-world applications requiring fast decision-making.
@@ 5 Extrinsic Evaluation
We develop RIGRec, a novel recommendation model that leverages the rich semantic relationships captured in our intention knowledge graph. The model employs graph convolution to derive informative item representations:$\bm{E}^{l+1}=\bm{A}\bm{E}^{l},$ where $\bm{A}$ represents the adjacency matrix of graph $G$, and $\bm{E}\in\mathbb{R}^{N\times d}$ stores the $d$-dimensional embeddings for all $N$ items. We implement lightweight convolution operations and sum pooling to enhance computational efficiency. After $L$ convolution layers, the resulting representations $\bm{E}^{L}$ capture both item characteristics and their intention-based relationships. For session modeling, we adopt SASRec’s self-attention architecture, which effectively aggregates item representations within a session to model user preferences. This creates an end-to-end recommendation system that seamlessly integrates intention-level reasoning with sequential pattern recognition, enabling more contextually appropriate recommendations.
@@
We employ two standard metrics in the field of recommender systems, including Recall at a cutoff top $k$ (Recall@$k$) and Normalized Discounted Cumulative Gain at a cutoff top $k$ (NDCG@$k$). We rank the ground-truth item alongside all candidates to ensure an unbiased evaluation rather than adopting the negative sampling strategy. We report the averaged metrics over 5 runs with the commonly utilized $k\in\{5,10,20,50,100\}$. To ensure unbiased evaluation, we rank the ground-truth item alongside all candidate items rather than using negative sampling. Results represent averages across five independent runs, with statistical significance determined through unpaired t-tests (p < 0.05). The implementation details are in the Appendix A.5.
@@ Appendix D Evaluating Diversity of Intention
$\text{Diversity}(D,n)=\frac{\#\text{ unique }n\text{-grams in }D^{\oplus}}{\#\text{ n-grams in }D^{\oplus}}$ (1)

### v3

@@ Abstract
Understanding user intentions is challenging for online platforms. Recent work on intention knowledge graphs addresses this, but often lacks focus on connecting intentions, which is crucial for modeling user behavior and predicting future actions. This paper introduces a framework to automatically generate an intention knowledge graph, capturing connections between user intentions. Using the Amazon m2 dataset, we construct an intention graph with 351 million edges, demonstrating high plausibility and acceptance. Our model effectively predicts new session intentions and enhances product recommendations, outperforming previous state-of-the-art methods and showcasing the approach’s practical utility. [^1] https://github.com/HKUST-KnowComp/RelationalIntentionGraph
@@ 2 Related Work
Meanwhile, knowledge graphs have also demonstrate its potential in solving various types of complex logical questions Bai et al. (2022, 2023c, 2023b, 2023a, 2024b); Zheng et al. (2025) in e-commerce domain Bai et al. (2024a). Automatic knowledge graph construction with intentions can also provide resource to the construction of neural graph databases Besta et al. (2022); Ren et al. (2023); Hu et al. (2024); Bai et al. (2025); Hu et al. (2025).
@@ 4.2 Automatic Evaluation
This study evaluates the task of linking unseen sessions to an existing intention graph. The dataset, comprising all session-intention edges, was split into training, validation, and test sets (8:1:1) based on sessions, with each session linked to 2-4 positive intentions. For evaluation, we applied random negative sampling to generate 30 candidate intentions per session, ranking the positive intentions among them. We compared large language models (Mistral-7B-Instruct-v0.1, Meta-Llama-3-8B-Instruct, and flan-t5-large) to our RIG model. These models are open-sourced and has the licenses that allow academic use. Baselines use perplexity scores derived via proper prompting to rank intentions, while RIG employs an embedding-based approach. Specifically, we generate embeddings for sessions and intentions using a sentence model and use SASRec as a session encoder to compute session representations. As shown in Table 7, our RIG model achieves competitive accuracy (e.g., best Hit@10) while significantly outperforming LLMs in inference speed (3.01ms vs. 2,021–4,102ms), demonstrating its practicality for real-world applications requiring fast decision-making.
@@ 5 Extrinsic Evaluation
[…]
@@
We employ two standard metrics in the field of recommender systems, including Recall at a cutoff top $k$ (Recall@k@k) and Normalized Discounted Cumulative Gain at a cutoff top $k$ (NDCG@k@k). We rank the ground-truth item alongside all candidates to ensure an unbiased evaluation rather than adopting the negative sampling strategy. We report the averaged metrics over 5 runs with the commonly utilized $k\in\{5,10,20,50,100\}$. To ensure unbiased evaluation, we rank the ground-truth item alongside all candidate items rather than using negative sampling. Results represent averages across five independent runs, with statistical significance determined through unpaired t-tests (p < 0.05). The implementation details are in the Appendix A.5.
@@ Appendix D Evaluating Diversity of Intention
[…]
@@ F.4 Analysis of Intention Relations
Appendix G Extended Ablation Study on LLM Backbones for Intention Generation
To assess the robustness of our framework and address concerns regarding the reliance on proprietary models (specifically GPT-3.5), we conducted an ablation study comparing different Large Language Models (LLMs) for the intention generation phase. We compared the current backbone (GPT-3.5) against two state-of-the-art open-weight models: Llama3-8B-Instruct and Mistral-7B-Instruct.
We evaluated the generated intentions based on Plausibility (whether the intention makes sense in the context) and Typicality (how common the intention is), alongside the inference cost.
Model Plausibility Typicality Cost (per 1K tokens)
GPT-3.5 (Ours) 0.955 0.667 $\sim$\$0.0015
Llama3-8B 0.948 0.652 Free (Self-hosted)
Mistral-7B 0.931 0.638 Free (Self-hosted)
Table 16: Performance and Cost Comparison of LLM Backbones. Our framework maintains high performance even with open-weight models.
Analysis.
As shown in Table 16, while GPT-3.5 achieves the highest performance, the gap between it and open-weight models is marginal. Llama3-8B trails by only 0.7% in plausibility and 1.5% in typicality. This demonstrates that our IGC-RC (Intention Generation, Conceptualization, and Relation Classification) framework is model-agnostic and can be effectively deployed using open-source models in resource-constrained environments or where data privacy requires local processing.
Appendix H Sensitivity Analysis of Vera Score Thresholds
In the construction of the Relational Intention Graph (RIG), we utilized a Vera score threshold to filter low-quality edges. To validate our choice of threshold and analyze the trade-off between graph scale and edge acceptance quality, we performed a sensitivity analysis using thresholds of 0.85, 0.90 (our selected value), and 0.95.
Threshold Acceptance Rate Number of Edges
0.85 79.8% 412.0 M
0.90 (Selected) 81.2% 341.6 M
0.95 82.4% 218.0 M
Table 17: Impact of Vera Score Thresholds on Graph Scale and Quality. The selected threshold (0.90) offers the optimal trade-off.
Analysis.
The results in Table 17 highlight the rationale behind our selection. A threshold of 0.95 significantly reduces the graph density (dropping nearly 123 million edges) for a marginal gain in acceptance rate (1.2%). Conversely, lowering the threshold to 0.85 introduces more noise. The 0.90 threshold offers the optimal balance, maintaining a high acceptance rate ($80\%$) while preserving the massive scale necessary for training robust recommendation models.
[…]
Appendix I Inference Efficiency and Practicality Analysis
A key advantage of our pre-computed Knowledge Graph approach (RIG) over direct LLM inference for recommendations is speed. We compared the inference latency of our RIG-enhanced recommendation model against a direct LLM-based recommender (using Mistral-7B).
Method MRR Hit@10 Inference Time Speedup
[…]
Table 18: Efficiency Comparison (Inference Time vs. Performance). RIG achieves comparable accuracy with drastically lower latency.